# Price predictability at ultra-high frequency: Entropy-based randomness test

### v2

@@ 3 Dataset
Asset
Ticker
Mean price
Standard deviation of price
Daily trading volume
Daily number of transactions
Average time between transactions
Apple Inc.
AAPL
153.47
0.93
12,184,032
136,136
0.165
Microsoft Corporation
MSFT
251.78
1.37
4,529,093
84,342
0.269
Tesla Inc.
TSLA
388.02
3.81
8,686,354
178,704
0.127
Intel Corporation
INTC
30.15
0.20
7,055,642
38,255
0.595
Eli Lilly and Company
LLY
327.33
1.73
370,050
11,404
2.086
Snap Inc.
SNAP
10.67
0.14
4,967,779
18,521
1.358
Ford Motor Company
F
13.93
0.10
4,468,175
12,954
1.815
Carnival Corporation & plc
CCL
9.24
0.12
5,874,376
15,372
1.518
SPDR S&P 500 ETF
SPY
390.52
1.56
9,136,137
95,181
0.246
@@
Considering each occurred transaction, we work with data in transaction time [38]. Discretization is made by distinguishing between positive and negative returns: 00 corresponds to price decreasing, $1$ corresponds to price increasing. Thus, alphabet $A$ is $\{0,1\}$ and a symbolic sequence is obtained according to binary discretization from Eq. 3.
$\begin{split}s^{(2)}_{t}=\begin{cases}0,r_{t}<0,\\ 1,r_{t}>0\end{cases}\end{split}$ (3)
@@ 4.1 Modeling ultra-high frequency data
We first apply the proposed method for testing predictability on simulated data. Executed orders and trade signs are generated using models that replicate order splitting, herd or chartist behavior, and mean-reverting process [39, 40, 41]. Positive autocorrelation in order flow was previously observed in [18] and [39], where authors demonstrated that autocorrelation diminishes with increasing aggregation by the number of orders. The $\lambda$ model proposed in [40] reproduces the property of the positive autocorrelation decreasing with the larger time lags. The $\lambda$ model introduces a fluctuating number of hidden orders that are divided into equal pieces and submitted gradually. We simulate 80 sequences of signs according to the $\lambda$ model with the length $10^{5}$ and parameters calibrated in the paper [^3] The number of hidden orders is $N=21$, the parameter of Pareto distribution describing volumes is $\alpha=1.63$, and the probability of a new hidden order is $\lambda=0.38$.. We present the fraction of predictable sequences for time lags from 1 to 50 in Figure 0(a) for the $\lambda$ model and two models described below. For small time lags, all sequences are predictable, then the fraction decreases but not monotonically.
@@
Lastly, we consider the trade superposition (TS) model proposed in [39]. This model posits that each price change results from previous trades, with a specific trade’s impact defined by a propagator that diminishes over time [^5] The model’s parameters, sourced from the paper, specify the propagator as $\frac{2.8\times 10^{-3}}{(l+20)^{0.42}}$. The logarithm of volumes follows a normal distribution $N(5.5,1.8)$ and the noise terms have a standard deviation of $0.01$.. Figures 0(a) and 0(b) display the fractions of predictable sequences for order signs and price returns, respectively, based on the TS model.
@@ 4.2 Analysis of real data
Our interest lies in identifying additional factors contributing to predictability beyond symbol repetition due to long memory. Consequently, we proceed by analyzing aggregated high-frequency data. We use transaction time, that is, we aggregate by a number of transactions. We record the last available price for each time step. We examine the predictability of assets over several months and with different aggregation levels. We plot the fraction of days which is classified as predictable for the four months under consideration in Fig. 5.
@@ 4.3 Statistics of predictable time intervals
Bouchaud et al. [45] investigated statistical properties of the limit order books of several stocks. In particular, the authors stated that the distribution of changes in limit order prices exhibits a power-law tail. Engle and Russell [46] noted that the longest duration between transactions appeared in the middle of trading days. According to the authors, clustering of transactions appears because of the size of bid-ask spread and gathering of informed traders. U-shapes of frequencies of large trades, small trades, and market orders were also discovered in [47]. In another paper by Engle [48], he showed that intraday volatility has the similar pattern and attains its minimum in the middle of a trading day. Moreover, significant coefficients of ARMA(1,1) model were detected. Highly dependent microstructure noise was stated by [49]. According to [50], changes in stock prices between transactions are associated with trading volumes. Some stylized facts including fat tails of price returns, volatility clustering, and leverage effect were discussed in [51]. For a more comprehensive understanding and detailed discussions regarding market and limit orders, we refer to works [37, 47, 52].
We show that some assets exhibit unpredictable prices for a part of days. Now, we consider different parameters of price returns time series and check if there is a dependence between them and predictability. The list of parameters follows.
  • First, we calculate the amount of non-zero returns that is the length of the symbolized sequence and the fraction of 0-returns.
  • Then, we record lengths of blocks, $k$. The amount of price returns used as past observations to test the independence is $k-1$.
@@
  • We are interested in autocorrelation of non-zero returns as well as in autocorrelation of their magnitude values. The autocorrelation of magnitudes is a proxy for volatility clustering.
  • Then, we consider the distribution of price returns. We fit empirical price returns distribution [^6] Here, we use scipy.stats.t.fit in Python. and record the degrees of freedom $\nu$, scale, and shift parameters. The smaller value of $\nu$, the fatter tails of price returns.
@@
We conduct a statistical test to examine the differences in mean values between predictable and unpredictable days. A p-value below 0.05 indicates a significant difference with a 95% confidence level. In the case of AAPL stock, all 80 days are determined as predictable using the NP-statistics, while 79 predictable days are detected via entropy estimation. To collect days where the hypothesis of unpredictability can not be rejected, we aggregate price returns. We start from the aggregation level $a=15$. Mean values of the specified parameters for AAPL stock are presented in Table 2, while Table 3 displays the parameter comparison between predictable and unpredictable days for all assets.
@@
parameter
mean for predictable days
mean for not predictable days
p-value
Sample size (number of days)
35
45
number of non-zero returns
5225
4370
0.009**
fraction of 0-returns
0.142
0.138
0.320
$k$
6
5.889
0.118
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$
3.084
2.635
3.152×$\times 10^{-5}$**
$|\hat{p}(1)-\hat{p}(0)|$
0.051
0.050
0.874
magnitude of daily log-price increment
0.017
0.012
0.093
mean price returns
2.185×$\times 10^{-7}$
-6.673×$\times 10^{-7}$
0.315
magnitude of autocorrelation of non-zero returns
0.044
0.027
0.002**
magnitude of autocorrelation of absolute values
0.115
0.088
0.007**
$\nu$ of t-distribution
2.220
2.694
0.31
scale of t-distribution
1.367×$\times 10^{-4}$
1.464×$\times 10^{-4}$
0.195
magnitude of shift of t-distribution
9.082×$\times 10^{-6}$
9.991×$\times 10^{-6}$
0.519
daily volume
13,367,901
11,257,199
0.026*
fraction of jumps
1.934×$\times 10^{-4}$
1.367×$\times 10^{-4}$
0.262
@@
parameter
AAPL MSFT TSLA INTC LLY SNAP F CCL SPY
aggregation level
10 30 10 25 2 5 1 1 1 5
number of predictable days
35 3 26 35 44 44 53 41 38 29
number of non-zero returns
fraction of 0-returns
$k$
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$
magnitude of daily log-price increment
mean price returns
magnitude of autocorrelation of non-zero returns
magnitude of autocorrelation of absolute values
$\nu$ of t-distribution
scale of t-distribution
daily volume
fraction of jumps
@@ 4.4 Predictability of pairs of signs
For the majority of considered assets, we find that predictability is associated with a high frequency of blocks with the same price direction. However, for three stocks, SNAP, F, CCL, we fail to detect the high frequency of blocks $0\dots 0$ and $1\dots 1$ without aggregation of prices. Here, we consider the three stocks to test if a price direction depends on the previous recorded decrease or increase. In other words, we test $\mathcal{H}_{0}$ setting $k=2$ and estimate $p(00)+p(11)$ for predictable and not predictable days. A summary table is Table 4. For all three stocks we detect a significant difference in $\hat{p}(00)+\hat{p}(11)$ for the two groups of days. However, the sum of two probabilities is less than $0.5$ for predictable days. That is, the probability of changing price direction is significantly high for predictable days. Therefore, predictable days are more likely to have a pattern of changing symbols indicating an increase or decrease in price.
@@
Stock
N. of predictable days
$\hat{p}(00)+\hat{p}(11)$ for predictable days
$\hat{p}(00)+\hat{p}(11)$ for unpredictable days
p-value for difference in mean values
SNAP
36
0.469
0.501
0.0014
F
25
0.460
0.493
2.671×$\times 10^{-10}$
CCL
27
0.478
0.5
0.044
@@ 4.5 Localization of predictable intervals
There are three days for AAPL stock where predictability is detected when the data is aggregated. For all three days predictability is detected only for daily time intervals ($S=1$). For the MSFT stock, there are three days where predictability is detected only for the part of the day ($S>1$). Regularity patterns are detected at the end of the day for August 3 and October 26. For August 5, predictability is presented for the first half of the day. It disappears at the next subsequent non-overlapping interval. Regarding TSLA stock, detectable predictability was identified in both halves of the day on August 5, indicating the potential to reduce the sequence length for predictability analysis by half and still observe significant patterns throughout the day. For the stock SNAP, we detect days when predictable time intervals occur several times during a day. We notice that some predictable intervals cluster together and highlight such intervals in the Table 8. For instance, we observe 7 predictable intervals going in a row on October 21 and 24. For 13 out of 30 days where $S_{max}>1$ for the Ford stock, the predictability can be found in the second half of a day. Similarly, there are 14 days in the sample of 31 days with $S_{max}>1$ where the predictability is detected only in the second part of a trading day for CCL. For each predictable day of the ETF SPY, there is only one from $S$ intervals where predictability is found. Predictability disappears from the time interval to the next subsequent non-overlapping interval. For instance, predictability was noticeable in the middle of the trading day on August 17.
@@ 5 Discussion and Conclusions
  • First, we estimate the Shannon entropy using empirical frequencies of blocks of symbols as suggested in [21]. Using empirical frequencies obtained by rolling a window with a certain length, we the calculate the NP-statistics.
@@
  • The statistics is scaled and shifted KL-divergence [36] that measures difference between empirical probabilities and theoretical probabilities under the null hypothesis of unpredictability.
@@
We notice that predictable days of AAPL, MSFT, INTC, LLY, F, CCL, and SPY are characterized by larger trading volumes and the larger amount of non-zero price changes in comparison with not predictable days. This distinction in characteristics suggests a correlation between predictability and trading activity, particularly in response to market news events. This assumption aligns with existing research demonstrating the influence of news on stock prices. Stock prices react to announcements about stock dividends and splits [59]. Weekly price returns react to attention to news and their tone [60]. Public news affect monthly price returns [61].
@@ Appendix A Asymptotic distribution of the Neyman-Pearson statistics
We outline the proof of Lemma 1. The lemma gives the asymptotic distribution of the statistics $D$ in Eq. 2 based on empirical frequencies of blocks of symbols of a sequence. If symbols of sequence $X=\{x_{t}\}_{t=1}^{n}$ are independently distributed, then the sequence $\bar{X}$ of blocks $\bar{x}_{t}=\{x_{t},x_{t+1},\dots,x_{t+k-2}\}$ follows the first order Markov process. The transition matrix of the Markov chain has dimension $M\times M$, where $M=s^{k-1}$. However, there are only $s$ non-zero probabilities whose sum is equal to one in each row of the matrix. Thus, the transition matrix can be converted into $M\times s$ matrix with only positive entries. This transformation is feasible since an output $\bar{x}_{t}$ in sequence $\bar{X}$ differs from $\bar{x}_{t-1}$ solely by a single new symbol $x_{t+k-2}\in A$.

### v3

@@ 3 Dataset
Asset Ticker Mean price Standard deviation of price Daily trading volume Daily number of transactions Average time between transactions
Apple Inc. AAPL 153.47 0.93 12,184,032 136,136 0.165
Microsoft Corporation MSFT 251.78 1.37 4,529,093 84,342 0.269
Tesla Inc. TSLA 388.02 3.81 8,686,354 178,704 0.127
Intel Corporation INTC 30.15 0.20 7,055,642 38,255 0.595
Eli Lilly and Company LLY 327.33 1.73 370,050 11,404 2.086
Snap Inc. SNAP 10.67 0.14 4,967,779 18,521 1.358
Ford Motor Company F 13.93 0.10 4,468,175 12,954 1.815
Carnival Corporation & plc CCL 9.24 0.12 5,874,376 15,372 1.518
SPDR S&P 500 ETF SPY 390.52 1.56 9,136,137 95,181 0.246
@@
Considering each occurred transaction, we work with data in transaction time [36]. Discretization is made by distinguishing between positive and negative returns: 00 corresponds to price decreasing, $1$ corresponds to price increasing. Thus, alphabet $A$ is $\{0,1\}$ and a symbolic sequence is obtained according to binary discretization from Equation 6:
$\begin{split}s^{(2)}_{t}=\begin{cases}0,r_{t}<0,\\ 1,r_{t}>0,\end{cases}\end{split}$ (6)
@@ 4.1 Modeling ultra-high frequency data
We first apply the proposed method for testing predictability on simulated data. We select models of limit orders exhibiting predictability patterns. We study how the predictability level changes with the aggregation of prices. Executed orders and trade signs are generated using models that replicate order splitting, herd or chartist behavior, and mean-reverting process [17, 37, 38]. According to Kyle’s model [39] including informed traders, a price return linearly depends on the trading volume. In this model, the volume traded by an informed trader depends on the difference between current price and its future fundamental value. According to a behavioral stock market model proposed in [22], prices may deviate from their fundamental value due to the behavior of sentiment traders, who buy stocks in rising markets. Positive autocorrelation in order flow was previously observed in [16] and [17], where the authors demonstrated that autocorrelation diminishes with increasing aggregation by the number of orders. The $\lambda$ model proposed in [37] reproduces the property of the positive autocorrelation decreasing with the larger time lags. The $\lambda$ model introduces a fluctuating number of hidden orders that are divided into equal pieces and submitted gradually. We simulate 80 sequences of signs according to the $\lambda$ model with the length $10^{5}$ and parameters calibrated in the paper [^2] The number of hidden orders is $N=21$, the parameter of Pareto distribution describing volumes is $\alpha=1.63$, and the probability of a new hidden order is $\lambda=0.38$.. We present the fraction of predictable sequences for time lags from 1 to 50 in Figure 1(a) for the $\lambda$ model and two models described below. For small time lags, all sequences are predictable, then the fraction decreases but not monotonically.
@@
Lastly, we consider the trade superposition (TS) model proposed in [17]. This model posits that each price change results from previous trades, with a specific trade’s impact defined by a propagator that diminishes over time [^4] The model’s parameters, sourced from the paper, specify the propagator as $\frac{2.8\times 10^{-3}}{(l+20)^{0.42}}$. The logarithm of volumes follows a normal distribution $N(5.5,1.8)$ and the noise terms have a standard deviation of $0.01$.. Figures 1(a) and 1(b) display the fractions of predictable sequences for order signs and price returns, respectively, based on the TS model. As we will see later in Figure 5 for real data analysis, the predictability of aggregated prices of frequently traded stocks is similar to the predictability of the TS model.
@@ 4.2 Analysis of real data
Our interest lies in identifying additional factors contributing to predictability beyond symbol repetition due to price persistence. Consequently, we proceed by analyzing aggregated high-frequency data. We use transaction time, that is, we aggregate by a number of transactions. We record the last available price for each time step. We examine the predictability of assets over several months and with different aggregation levels. We plot the fraction of days which is classified as predictable for the four months under consideration in Fig. 5.
@@ 4.3 Statistics of predictable time intervals
We show that some assets exhibit unpredictable prices for a part of days. In this section, we explore what price characteristics distinguish days with predictability from others. We first review stylized facts about ultra-high frequency data.
Lillo and Farmer [16] conducted a statistical test, concluding that both the signs of market orders and executed limit orders exhibit long-memory processes. They attributed this long memory of order signs to news arrivals and order splitting, which is offset by fluctuations in market liquidity. Therefore, the long memory does not contradict efficiency of markets when prices incorporate all available information about future values [43]. Bouchaud et al. [44] investigated statistical properties of the limit order books of several stocks. In particular, the authors stated that the distribution of changes in limit order prices exhibits a power-law tail. Engle and Russell [45] noted that the longest duration between transactions appeared in the middle of trading days. According to the authors, clustering of transactions appears because of the size of bid-ask spread and gathering of informed traders. U-shapes of frequencies of large trades, small trades, and market orders were also discovered in [46]. In another paper by Engle [15], he showed that intraday volatility has the similar pattern and attains its minimum in the middle of a trading day. Moreover, significant coefficients of ARMA(1,1) model were detected. Highly dependent microstructure noise was stated by [47]. According to [48], changes in stock prices between transactions are associated with trading volumes. Some stylized facts including fat tails of price returns, volatility clustering, and leverage effect were discussed in [49]. For a more comprehensive understanding and detailed discussions regarding market and limit orders, we refer to works [35, 46, 50].
Now, we consider different parameters of price returns time series and check if there is a dependence between them and predictability. The list of parameters follows.
  • First, we calculate the amount of non-zero returns that is the length of the symbolized sequence $n$ and the fraction of 0-returns. Then, we record lengths of blocks, $k=\left[0.5\log_{s}{n}\right]$. The amount of price returns used as past observations to test the independence is $k-1$.
@@
  • We are interested in autocorrelation with lag 1 of non-zero returns as well as in autocorrelation of their magnitude values. The autocorrelation of magnitudes is a proxy for volatility clustering.
  • Then, we consider the distribution of price returns. We fit a Student’s distribution of returns distribution and record the degrees of freedom $\nu$, scale, and shift parameters [^5] Here, we use scipy.stats.t.fit in Python.. The smaller value of $\nu$, the fatter tails of price returns.
@@
We conduct a statistical test to examine the differences in mean values between predictable and unpredictable days. A p-value below 0.05 indicates a significant difference with a 95% confidence level. We aggregate prices to collect both sets of predictable and not predictable days. The difference in characteristics may vary depending on the level of aggregation. For example, the persistence of the return sign is more typical for higher frequencies.
In the case of AAPL stock, all 80 days are determined as predictable using the NP-statistics, while 79 predictable days are detected via entropy estimation. To collect days where the hypothesis of unpredictability can not be rejected, we aggregate price returns. We start from the aggregation level $a=15$. Setting the aggregation level to 15 means that we consider only every 15th transaction and use them to build a symbolic series of price increases or decreases. The level of aggregation is chosen to balance between predictable and unpredictable days. Mean values of the specified parameters for AAPL stock are presented in Table 2, while Table 3 displays the parameter comparison between predictable and unpredictable days for all assets.
@@
parameter mean for predictable days mean for not predictable days p-value
Sample size (number of days) 35 45
number of non-zero returns 5225 4370 0.009**
fraction of 0-returns 0.142 0.138 0.320
$k$ 6 5.889 0.118
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$ 3.084 2.635 3.152×$\times 10^{-5}$**
$|\hat{p}(1)-\hat{p}(0)|$ 0.051 0.050 0.874
magnitude of daily log-price increment 0.017 0.012 0.093
mean price returns 2.185×$\times 10^{-7}$6.673×$\times 10^{-7}$ 0.315
magnitude of autocorrelation of non-zero returns 0.044 0.027 0.002**
magnitude of autocorrelation of absolute values 0.115 0.088 0.007**
$\nu$ of t-distribution 2.220 2.694 0.31
scale of t-distribution 1.367×$\times 10^{-4}$ 1.464×$\times 10^{-4}$ 0.195
magnitude of shift of t-distribution 9.082×$\times 10^{-6}$ 9.991×$\times 10^{-6}$ 0.519
daily volume 13,367,901 11,257,199 0.026*
fraction of jumps 1.934×$\times 10^{-4}$ 1.367×$\times 10^{-4}$ 0.262
@@
parameter AAPL MSFT TSLA INTC LLY SNAP F CCL SPY
aggregation level 10 30 10 25 2 5 1 1 1 5
number of predictable days 35 3 26 35 44 44 53 41 38 29
number of non-zero returns $>$** $>$** $>$* $>$** $>$** $>$** $>$** $>$* $>$*
fraction of 0-returns $>$* $>$** $>$**
$k$ $>$** $>$* $>$** $>$*
$(\hat{p}(0\dots 0)+\hat{p}(1\dots 1))2^{k}$ $>$** $>$** $>$** $>$** $>$** $>$**
magnitude of daily log-price increment $>$**
mean price returns $>$* $>$*
magnitude of autocorrelation of non-zero returns $>$** $>$** $>$** $>$** $>$** $>$** $>$** $>$**
magnitude of autocorrelation of absolute values $>$** $<$** $>$** $>$* $>$**
$\nu$ of t-distribution $<$* $>$* $>$** $<$*
scale of t-distribution $<$* $<$* $<$*
daily volume $>$* $>$** $>$* $>$** $>$** $>$** $>$** $>$*
fraction of jumps $<$** $>$*
@@ 4.4 Predictability of pairs of signs
Here, we explore the predictability of pairs of price changes, that is the randomness of future price movement observing the last price change. We set the user-defined parameter $k$ to be equal to 2. For the majority of considered assets, we find that predictability is associated with a high frequency of blocks with the same price direction. However, for three stocks, SNAP, F, CCL, we fail to detect the high frequency of blocks $0\dots 0$ and $1\dots 1$ without aggregation of prices. Here, we consider the three stocks to test if a price direction depends on the previous recorded decrease or increase. In other words, we test $\mathcal{H}_{0}$ setting $k=2$ and estimate $p(00)+p(11)$ for predictable and not predictable days. A summary table is Table 4. For all three stocks we detect a significant difference in $\hat{p}(00)+\hat{p}(11)$ for the two groups of days. However, the sum of two probabilities is less than $0.5$ for predictable days. That is, the probability of changing price direction is significantly high for predictable days. Therefore, predictable days are more likely to have a pattern of changing symbols indicating an increase or decrease in price.
@@
Stock N. of predictable days $\hat{p}(00)+\hat{p}(11)$ for predictable days $\hat{p}(00)+\hat{p}(11)$ for unpredictable days p-value for difference in mean values
SNAP 36 0.469 0.501 0.0014
F 25 0.460 0.493 2.671×$\times 10^{-10}$
CCL 27 0.478 0.5 0.044
@@ 4.5 Localization of predictable intervals
There are three days for AAPL stock where predictability is detected when the data is aggregated. For all three days predictability is detected only for daily time intervals ($S=1$). For the MSFT stock, there are three days where predictability is detected only for the part of the day ($S>1$). Regularity patterns are detected at the end of the day for August 3 and October 26. For August 5, predictability is presented for the first half of the day. It disappears at the next subsequent non-overlapping interval. For each predictable day of the ETF SPY, there is only one from $S$ intervals where predictability is found. Predictability disappears from the time interval to the next subsequent non-overlapping interval. For instance, predictability was noticeable in the middle of the trading day on August 17. Regarding TSLA stock, detectable predictability was identified in both halves of the day on August 5, indicating the potential to reduce the sequence length for predictability analysis by half and still observe significant patterns throughout the day. For 13 out of 30 days where $S_{max}>1$ for the Ford stock, the predictability can be found in the second half of a day. Similarly, there are 14 days in the sample of 31 days with $S_{max}>1$ where the predictability is detected only in the second part of a trading day for CCL.
For the stock SNAP, we detect days when predictable time intervals occur several times during a day. We notice that some predictable intervals cluster together and highlight such intervals in the Table 8. For instance, we observe 7 predictable intervals going in a row on October 21 and 24. However, we do not find evidence that predictive intervals follow each other for the other stocks considered.
@@ 5 Discussion and Conclusions
  • First, we estimate the Shannon entropy using empirical frequencies of blocks of symbols as suggested in [19]. Using empirical frequencies obtained by rolling a window with a certain length, we calculate the NP-statistics.
@@
  • The statistics is a scaled KL-divergence [34] that measures difference between empirical probabilities and theoretical probabilities under the null hypothesis of unpredictability.
@@
We notice that predictable days of AAPL, MSFT, INTC, LLY, F, CCL, and SPY are characterized by larger trading volumes and the larger amount of non-zero price changes in comparison with not predictable days. The dependence between price changes and trading volumes was described in the paper of Kyle [39]. The distinction in the volumes and price changes suggests a correlation between predictability and trading activity, particularly in response to market news events. This assumption aligns with existing research demonstrating the influence of news on stock prices. Stock prices react to announcements about stock dividends and splits [58]. Weekly price returns react to attention to news and their tone [59]. Public news affect monthly price returns [60].
@@ Appendix A Asymptotic distribution of the Neyman-Pearson statistics
We outline the proof of Lemma 2. The lemma gives the asymptotic distribution of the statistics $D$ in Equation 3 based on empirical frequencies of blocks of symbols of a sequence. If symbols of sequence $X=\{x_{t}\}_{t=1}^{n}$ are independently distributed, then the sequence $\bar{X}$ of blocks $\bar{x}_{t}=\{x_{t},x_{t+1},\dots,x_{t+k-2}\}$ follows the first order Markov process. The transition matrix of the Markov chain has dimension $M\times M$, where $M=s^{k-1}$. However, there are only $s$ non-zero probabilities whose sum is equal to one in each row of the matrix. Thus, the transition matrix can be converted into $M\times s$ matrix with only positive entries. This transformation is feasible since an output $\bar{x}_{t}$ in sequence $\bar{X}$ differs from $\bar{x}_{t-1}$ solely by a single new symbol $x_{t+k-2}\in A$.